OpenStudy (anonymous):

Fred has 10 coins in a book bag. He has 4 quarters, 1 penny, and 5 dimes. If Fred reaches into his book bag and draws a coin without looking, what is the probability that he will draw a quarter?

OpenStudy (ddcamp):

4(favorable outcomes)/10(possible outcomes) 4/10= 2/5 = 0.4 = 40%

OpenStudy (anonymous):

thanks